How many possible samples of size n = 9 can be drawn from a population of size 12?


1
Expert's answer
2022-04-05T02:27:58-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }=\\
=\begin{pmatrix}
 12\\
 9
\end{pmatrix}=\cfrac{12! } {9! \cdot3! }=\\
=\cfrac{10\cdot11\cdot12}{2\cdot3}=220 ."